OpenStudy (anonymous):

Express the following in simplest radical form. All variables represent positive real numbers. square 9x^9y^4

OpenStudy (anonymous):

I know \[\sqrt{9}= \sqrt{3 }\] but what about the 9 and 4

OpenStudy (anonymous):

\[\sqrt{9x^9y^4}\]

ganeshie8 (ganeshie8):

\(\sqrt{9} \ne \sqrt{3}\)

ganeshie8 (ganeshie8):

thats incorrect ok :) i think u mean this : \(\sqrt{9} = 3\)

OpenStudy (anonymous):

That what I meant, so how to get the exponent.

ganeshie8 (ganeshie8):

\(\sqrt{9x^9y^4} \) \(3 \sqrt{x^9y^4} \)

ganeshie8 (ganeshie8):

you wanto see if something else can come out of radical

ganeshie8 (ganeshie8):

write \(y^4 = (y^2)^2\)

ganeshie8 (ganeshie8):

and, write \(x^9 = x^8.x = (x^4)^2.x\)

ganeshie8 (ganeshie8):

see if u can simplify further